you inherit $70,000 and plan to invest the money in two stocks paying 4% and 10% annual interest, respectively. How much should be invested at each rate if the total interest earned for the year is to be $4000?
Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 10.00% per annum ------------ Amount invested = y
70000
Interest----- 4000.00

Part I 4.00% per annum ---x
Part II 10.00% per annum ---y
Total investment
x + 1 y= 70000 -------------1
Interest on both investments
4.00% x + 10.00% y= 4000
Multiply by 100
4 x + 10 y= 400000.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -280000.00
Add this to (2)
0 x 6 y= 120000
divide by 6
y = 20000
Part I 4.00% $ 50000
Part II 10.00% $ 20000

CHECK
50000 --------- 4.00% ------- 2000.00
20000 ------------- 10.00% ------- 2000.00
Total -------------------- 4000.00